# V2Xum-LLM: Cross-Modal Video Summarization with Temporal Prompt Instruction Tuning

### v2

@@ Introduction
This section introduces our unified cross-modal video summarization framework, V2Xum-LLaMA, which employs the interleaved video frames along with temporal and task prompts as input and converts videos into multimodal summaries. Figure 2 depicts the architecture of our framework, where video frames are first encoded into visual representations, then combined with temporal and task prompts to obtain the interleaved input tokens for the LLaMA-2 decoder. Using the LLaMA decoder, our system can generate both video-to-video and video-to-text summaries. In addition, we design a temporal-aware decoding training strategy to enhance the alignment of text and video summaries. To sum up, our approach enables modality control, producing either video, text summaries, or both, and enhances the model’s understanding of the video timeline for more relevant and accurate outputs.
@@
The V2V summarization $A^{v}$ is defined as a sequence of temporal tokens sequence, i.e., $A^{v}=\{v_{i}\}^{M}_{i=1}$, which is a subset of temporal prompts sequence $T$ and $M\leq L$. The V2T summarization is a summarized caption in a natural language format. In our implementation shown as Figure 2, the V2T summarization also contains frame referring temporal tokens:
@@
where the temporal token $t_{j}$ is bind with the words or sentences that consist of the summarization of visual content and can be further extracted from $A^{t}$ to get the V2V summarization $A^{v}$. This is called temporal-aware decoding. An example is shown in Fig. 2. When decoding, the temporal tokens can be decoded together with the text summaries and represent the temporal position where the described visual content occurred in the input video.
@@
Method Cross- Modal LLM- Based TSH- Free V2T V2V V2VT
@@ B-4 M R-L C F1 FCLIP Cross-FCLIP
We use CLIP ViT-L/14@336 as the vision encoder and Vicuna-v1.5-7B/13B as the text decoder. The vision adapter is pre-trained with image-text pairs from LCS-558K (Liu et al. 2023a) dataset. Each tokenized temporal prompt takes four tokens. To facilitate model training, we normalize the video length into 100 via downsampling. For each setting, we train the V2Xum-LLaMA for 5 epochs with a learning rate of 1e-4 on 8$\times$×NVIDIA A100 GPUs.
@@
We use the VideoXum dataset to evaluate our model’s capability to cross-model video summarization. The experimental results are shown in Table 2. It can be summarized that our proposed method outperforms all the baseline models. Specifically, in V2V summarization, V2Xum-LLaMA achieves a 8.1% higher F1-Score than VTSUM-BLIP, alongside significant improvements in both Spearman and Kendall correlation metrics. In addition, V2Xum-LLaMA-13B achieves higher evaluation scores on the V2V summarization task than V2Xum-LLaMA-7B. On the contrary, V2Xum-LLaMA-7B performs better on V2T summarization. We attribute this result to the increased complexity of the V2T summarization compared to V2V summarization. Additionally, the VideoXum training set comprises only 8,000 examples, a quantity insufficient for effectively training models with large language decoders.
We also evaluate various models on our newly proposed Instruct-V2Xum dataset, as detailed in Table 3. The results indicate that the models are well-adapted to the dataset, exhibiting sound performance. Moreover, V2Xum-LLaMA-13B outperforms V2Xum-LLaMA-7B in V2T summarization, which we believe is due to the larger language models benefiting from the increased volume of training data.
@@
We evaluate V2Xum-LLaMA on VideoXum, TVSum, SumMe, and Instruct-V2Xum datasets. The results are shown in Table 2 and Table 4. It can be concluded that the unified video summarization using the language decoders in V2Xum-LLaMA can effectively perform traditional V2V summarization tasks. V2Xum-LLaMA outperforms all the previous methods that relied on task-specific regression heads for generating video summaries. This result indicates that LLMs with temporal prompts are capable of performing fine-grained video temporal understanding. The results presented in Table 3 affirm the validity of our proposed V2Xum dataset. It demonstrates that the model can properly fit the data.
@@
A conventional approach to V2V summarization involves training a regression head to assign an importance score to each frame, ranking frames based on these scores, and selecting the top Kframes to evaluate performance using metrics like the F1 score or Kendall/Spearman correlation against the ground truth. However, as demand for cross-modal video summarization grows, this method is inadequate for multimodal video summarization. Recent studies have begun to explore the use of language models for generating temporal and spatial references in videos (Huang et al. 2023; Li et al. 2024; Ren et al. 2023), demonstrating the viability of using language models to generate the video interval indexes. In addition, leveraging large language models’ text decoders for both V2V and V2T summarization tasks is a logical step. Therefore, in this study, we investigate how to prompt large VLMs to understand the fine-grained video content along with temporal information, how to maximally leverage the powerful capability of content understanding and reasoning of large language models, and how to achieve task controllability via natural language instructions. To that end, we first propose the temporal prompt mechanism, and then design the visual encoder that takes the interleaved video frames, temporal, and language as input. Instead of using the regression head to perform V2V summarization, we unify different video summarization tasks into one language decoder.

### v3

@@ Introduction
This section introduces our unified cross-modal video summarization framework, V2Xum-LLaMA, which employs the interleaved video frames along with temporal and task prompts as input and converts videos into multimodal summaries. Figure˜2 depicts the architecture of our framework, where video frames are first encoded into visual representations, then combined with temporal and task prompts to obtain the interleaved input tokens for the LLaMA-2 decoder. Using the LLaMA decoder, our system can generate both video-to-video and video-to-text summaries. In addition, we design a temporal-aware decoding training strategy to enhance the alignment of text and video summaries. To sum up, our approach enables modality control, producing either video, text summaries, or both, and enhances the model’s understanding of the video timeline for more relevant and accurate outputs.
@@
The V2V summarization $A^{v}$ is defined as a sequence of temporal tokens sequence, i.e., $A^{v}=\{v_{i}\}^{M}_{i=1}$, which is a subset of temporal prompts sequence $T$ and $M\leq L$. The V2T summarization is a summarized caption in a natural language format. In our implementation shown as Figure˜2, the V2T summarization also contains frame referring temporal tokens:
@@
where the temporal token $t_{j}$ is bind with the words or sentences that consist of the summarization of visual content and can be further extracted from $A^{t}$ to get the V2V summarization $A^{v}$. This is called temporal-aware decoding. An example is shown in Fig.˜2. When decoding, the temporal tokens can be decoded together with the text summaries and represent the temporal position where the described visual content occurred in the input video.
@@
Method
Cross-
Modal
LLM-
Based
TSH-
Free
V2T V2V V2VT
@@ B-4 M R-L C F1 FCLIP Cross-FCLIP
[…]
@@
We use the VideoXum dataset to evaluate our model’s capability to cross-model video summarization. The experimental results are shown in Table˜2. It can be summarized that our proposed method outperforms all the baseline models. Specifically, in V2V summarization, V2Xum-LLaMA achieves a 8.1% higher F1-Score than VTSUM-BLIP, alongside significant improvements in both Spearman and Kendall correlation metrics. In addition, V2Xum-LLaMA-13B achieves higher evaluation scores on the V2V summarization task than V2Xum-LLaMA-7B. On the contrary, V2Xum-LLaMA-7B performs better on V2T summarization. We attribute this result to the increased complexity of the V2T summarization compared to V2V summarization. Additionally, the VideoXum training set comprises only 8,000 examples, a quantity insufficient for effectively training models with large language decoders.
We also evaluate various models on our newly proposed Instruct-V2Xum dataset, as detailed in Table˜3. The results indicate that the models are well-adapted to the dataset, exhibiting sound performance. Moreover, V2Xum-LLaMA-13B outperforms V2Xum-LLaMA-7B in V2T summarization, which we believe is due to the larger language models benefiting from the increased volume of training data.
@@
We evaluate V2Xum-LLaMA on VideoXum, TVSum, SumMe, and Instruct-V2Xum datasets. The results are shown in Table 2 and Table 4. It can be concluded that the unified video summarization using the language decoders in V2Xum-LLaMA can effectively perform traditional V2V summarization tasks. V2Xum-LLaMA outperforms all the previous methods that relied on task-specific regression heads for generating video summaries. This result indicates that LLMs with temporal prompts are capable of performing fine-grained video temporal understanding. The results presented in Table˜3 affirm the validity of our proposed V2Xum dataset. It demonstrates that the model can properly fit the data.
@@
A conventional approach to V2V summarization involves training a regression head to assign an importance score to each frame, ranking frames based on these scores, and selecting the top K% frames to evaluate performance using metrics like the F1 score or Kendall/Spearman correlation against the ground truth. However, as demand for cross-modal video summarization grows, this method is inadequate for multimodal video summarization. Recent studies have begun to explore the use of language models for generating temporal and spatial references in videos (Huang et al. 2023; Li et al. 2024; Ren et al. 2023), demonstrating the viability of using language models to generate the video interval indexes. In addition, leveraging large language models’ text decoders for both V2V and V2T summarization tasks is a logical step. Therefore, in this study, we investigate how to prompt large VLMs to understand the fine-grained video content along with temporal information, how to maximally leverage the powerful capability of content understanding and reasoning of large language models, and how to achieve task controllability via natural language instructions. To that end, we first propose the temporal prompt mechanism, and then design the visual encoder that takes the interleaved video frames, temporal, and language as input. Instead of using the regression head to perform V2V summarization, we unify different video summarization tasks into one language decoder.